OpenStudy (anonymous):

A recursive function is shown below: f(1) = 4 and f(n) = f(n - 1) - 8; n > 1 Which of the following lists the terms in the sequence defined by this recursive function? f(1) = 4 and f(n) = f(n - 1) - 8; n > 1 Which of the following lists the terms in the sequence defined by this recursive function? 4, -4, -12, -20, -28, ... 4, 4, 12, 20, 28, ... 4, 12, 20, 28, 36, ... 4, -12, -20, -28, -36, ... need help

OpenStudy (irishboy123):

if in doubt, why not list them out f(1) = 4 f(2) = f(2 - 1) - 8 = f(1) - 8 = -4 f(3) = ...... ??